### v1

# BenchPress: A Human-in-the-Loop Annotation System for Rapid Text-to-SQL Benchmark Curation

Abstract.
Large language models (LLMs) have been successfully applied to many tasks, including text-to-SQL generation. However, much of this work has focused on publicly available datasets, such as Fiben, Spider, and Bird. Our earlier work showed that LLMs are much less effective in querying large private enterprise data warehouses, releasing Beaver, the first private enterprise text-to-SQL benchmark. To create Beaver, we leveraged SQL logs, which are often readily available. However, manually annotating these logs to identify which natural language questions they answer is a daunting task. Asking database administrators, who are highly trained experts, to take on additional work to construct and validate corresponding natural language utterances is not only challenging but also quite costly.
To address this challenge, we introduce BenchPress, a human-in-the-loop system designed to accelerate the creation of domain-specific text-to-SQL benchmarks. Given a SQL query, BenchPress uses retrieval-augmented generation (RAG) and LLMs to propose multiple natural language descriptions. Human experts then select, rank, or edit these drafts to ensure accuracy and domain alignment. We evaluated BenchPress on annotated enterprise SQL logs, demonstrating that LLM-assisted annotation drastically reduces the time and effort required to create high-quality benchmarks. Our results show that combining human verification with LLM-generated suggestions enhances annotation accuracy, benchmark reliability, and model evaluation robustness. By streamlining the creation of custom benchmarks, BenchPress offers researchers and others a mechanism for assessing text-to-SQL models on a given domain-specific workload. BenchPress is freely available via our public GitHub repository [^1] https://github.com/fabian-wenz/enterprise-txt2sql and accessible for use on our website [^2] http://dsg-mcgraw.csail.mit.edu:5000/.
Key words and phrases:
Text-to-SQL, Benchmark Curation, Natural Language Interfaces, SQL Log Annotation, Large Language Models, Data Integration, Query Understanding, Database Usability, Enterprise Data, Human-in-the-Loop Annotation
{CCSXML}
<ccs2012> <concept> <concept_id>10002951.10002952.10003197</concept_id> <concept_desc>Information systems Query languages</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010178.10010179.10010186</concept_id> <concept_desc>Computing methodologies Language resources</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10002951.10003317</concept_id> <concept_desc>Information systems Information retrieval</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
1. Introduction
The adoption of large language models (LLMs) for text-to-SQL conversion has gained traction in enterprise settings, where databases are vast, but expert annotation resources are scarce. While academic research has produced powerful text-to-SQL models, enterprises face a critical problem: how well do these models perform on their data?
Public benchmarks like Spider Yu et al. (2018), BirdLi et al. (2023), and Fiben Sen et al. (2020) provide valuable testbeds for general-purpose text-to-SQL evaluation, but they fail to capture enterprise-specific challenges, such as:
  • Schema complexity and ambiguity: Enterprise databases are often heterogeneous, integrating tables from different systems with overlapping but inconsistent naming conventions.
  • Domain-specific terminology: Public datasets lack the specialized vocabulary and abbreviations used in finance, healthcare, IT, and other industries.
  • Privacy and security constraints: Unlike academic benchmarks, enterprise SQL logs cannot be publicly shared, making it difficult for organizations to benchmark models against real-world queries.
As a result, companies risk deploying models that fail on their data due to domain mismatch, leading to unreliable query generation and poor automation performance. While recent LLMs such as GPT-4o and fine-tuned LLaMA variants achieve impressive results on public datasets like Fiben, Spider, and Bird—to the point that the text-to-SQL task may appear nearly solved—, their execution accuracy  [^3] Execution accuracy measures whether the result of executing the predicted SQL query matches that of the gold SQL Yu et al. (2018); Li et al. (2023). drops sharply on enterprise datasets such as BeaverChen et al. (2024). Figure 1 visualizes this gap, showing a dramatic execution accuracy drop when the same models are evaluated on real-world enterprise queries. This discrepancy highlights the limitations of existing public benchmarks and the risk of overestimating model readiness for production use. To avoid deployment failures, enterprises must evaluate model performance under their own schemas, domain-specific terminology, and query patterns.
Figure 1. Execution accuracy of different LLMs on public benchmarks (Spider, Bird, Fiben) and the enterprise benchmark (Beaver). Since the best-performing model varies across datasets, the specific model achieving the highest accuracy is labeled above each teal-colored bar: Spider – miniSeek Seek AI Research Team (2023), Bird – askData Shkapenyuk et al. (2025), Fiben – Athena++ Sen et al. (2020), and BeaverChen et al. (2024) – contextModel.
To address this gap, we introduce BenchPress, a system designed to enable organizations to create their own text-to-SQL benchmarks quickly and efficiently. By combining LLM-generated suggestions with human validation, BenchPress produces accurate, domain-specific training data while reducing annotation effort. It offers a scalable and privacy-aware foundation for enterprise-grade text-to-SQL evaluation.
BenchPress operates in one primary way:
SQL-to-NL generation
Given a SQL query, BenchPress generates a natural language (NL) description, allowing domain experts to review and refine it, significantly reducing annotation effort.
By integrating LLM-generated suggestions with human validation, BenchPress accelerates annotation while maintaining accuracy, allowing enterprises to benchmark any text-to-SQL model against their own private datasets. We evaluated BenchPress on enterprise SQL logs, demonstrating its effectiveness in generating high-quality, domain-specific training data while reducing manual effort.
Beyond enabling benchmark creation, BenchPress has broader implications for enterprise AI adoption, allowing organizations to:
  • Assess model generalization to proprietary schemas before deployment.
  • Optimize fine-tuning strategies by identifying failure cases.
  • Improve interpretability of text-to-SQL models by systematically evaluating outputs.
By providing a scalable, adaptable, and privacy-aware solution for enterprise text-to-SQL benchmarking, BenchPress paves the way for more robust and domain-specific LLM evaluations.
2. Related Work
In this section, we summarize related work in benchmarks, annotation tools, LLM-based SQL generation, and enterprise adaptation.
Text-to-SQL Benchmarks
Public benchmarks such as Spider  Yu et al. (2018), Bird Li et al. (2023), and Fiben Sen et al. (2020) have driven progress in general-purpose text-to-SQL tasks. While these datasets capture diverse query types, they focus on clean, publicly available schemas. The Beaver benchmark Chen et al. (2024) introduces a combined corpus from academic and enterprise-inspired sources, including the network datasets and the datawarehouse from the Massachusetts Institute of Technology. Beaver demonstrates that LLMs struggle with more complex queries and heterogeneous schemas, motivating the need for domain-specific evaluation and adaptation tools. Unlike Beaver, which centers on model benchmarking, BenchPress focuses on enabling rapid benchmark creation through LLM-assisted human annotation.
LLMs for SQL Generation
Large language models such as Codex, GPT-4, and DeepSeek have shown strong results in translating natural language into SQL Chen et al. (2021); Ni et al. (2024). However, their performance degrades significantly in domain-specific or enterprise contexts. Maamari et al. Maamari et al. (2025) highlight this gap by studying LLMs in enterprise scenarios, finding that pre-trained models often fail due to domain-specific vocabulary and schema ambiguity. While their focus is on improving LLM robustness through better model training and prompting, our work complements this by tackling the data creation bottleneck—providing a system that enables enterprises to efficiently construct accurate, workload-specific training and evaluation data.
Annotation Systems and Tools
Several systems have explored semi-automated dataset creation. Andrejczuk et al. Andrejczuk et al. (2022) propose schema-aware data-to-text generation pipelines, and Xu et al. Xu et al. (2018) study SQL-to-text generation using encoder-decoder models. These approaches are primarily model-driven and assume public or synthetic data. In contrast, BenchPress supports human-in-the-loop workflows designed specifically for private enterprise logs, integrating LLM-generated suggestions with domain expert review.
Enterprise Data Challenges
Enterprise databases differ from academic benchmarks in terms of scale, schema complexity, and sensitivity. Prior work on federated training and schema linking Peng et al. (2022); Zhong et al. (2020) has explored solutions for isolated aspects, but few systems address the full annotation pipeline. BenchPress fills this gap by enabling benchmark creation that is both domain-adaptive and privacy-aware—without requiring public data release or costly in-house labeling from scratch.
BenchPress in Context
Unlike prior systems, BenchPress provides a practical toolkit for constructing domain-specific text-to-SQL benchmarks. It supports now mainly SQL-to-NL annotation for validation, semantic enrichment, and human verification, enabling scalable, robust benchmarking in enterprise settings. BenchPress provides a means to generate new, customized corpora—tailored to real-world workloads.
3. Enterprise SQL Logs: Challenges
In this research, we have worked with four text-to-SQL benchmarks: Spider, Bird, and Fiben as public datasets, and Beaver as a private, enterprise-oriented dataset. Beaver is based on SQL logs from enterprise databases across industries such as education, technology, and manufacturing. These logs span in total over 300 schemas and nearly 4000 queries, featuring complex, multi-source schemas with inconsistent naming conventions and semantic overlaps. For example, in the MIT data warehouse, a single natural language query can often be answered by multiple SQL queries due to the presence of materialized views and semantic ambiguity across tables.
Next we discuss the challenges of working with enterprise SQL logs in contrast to those from open-domain benchmarks:
Domain-Specific Terminology:
Enterprise SQL logs often contain specialized vocabulary that requires deep contextual understanding. For instance, terms like “J-term” (a one-month January term) are specific to the MIT academic calendar and may be incomprehensible to annotators or models without MIT-specific knowledge. Without domain expertise, LLMs frequently fail to map such terms to the correct database fields, reducing annotation and generation accuracy.
Query Complexity:
Queries in enterprise settings are substantially more complex than those in public benchmarks. They commonly include nested subqueries, aggregation dependencies, and recursive joins. A single query may aggregate data from 5–10 tables and use the result in a nested filter. LLMs, which are typically trained on simpler academic benchmarks, struggle to handle this structural depth. A comparative graph illustrating query complexity follows.
Privacy Constraints:
Enterprise data often contains sensitive or proprietary information, such as employee salaries or internal metrics. Due to strict privacy constraints, this data cannot be publicly released or used for model training. BenchPress addresses this challenge by enabling organizations to evaluate public models securely on private SQL logs—supporting informed model selection without compromising confidentiality.
Schema Ambiguity and Duplication:
Enterprise data warehouses often aggregate tables from various internal systems, leading to schema inconsistencies. It is common to find multiple tables with identically named columns like “user_id” that actually refer to different entities. Disambiguating such cases requires either careful schema documentation or intelligent annotation, both of which pose a challenge for LLM-based systems.
Data Sparsity and Imbalance:
Many enterprise datasets suffer from sparsity and imbalance. For example, the Intel data warehouse includes performance metrics from millions of devices, but many fields may be missing or unevenly distributed. LLMs trained on uniformly structured or synthetic data may struggle with these irregularities, resulting in biased performance on common patterns and failures on rare or incomplete cases.
In summary, enterprise SQL logs differ fundamentally from public benchmarks. The challenges span domain-specific terminology, complex query structures, privacy restrictions, ambiguous schemas, and sparse or imbalanced data. Addressing these issues demands tools like BenchPress that incorporate human-in-the-loop annotation, secure evaluation workflows, and domain-aware disambiguation mechanisms.
4. The BenchPress System
BenchPress is a human-in-the-loop system designed to accelerate the annotation of SQL logs for building high-quality text-to-SQL benchmarks. It enables domain experts to generate natural language descriptions for SQL queries more efficiently by combining retrieval-augmented generation (RAG), prompt-based LLM outputs, and iterative feedback refinement in a modular and interactive interface.
4.1. Workflow Overview
Figure 2. BenchPress workflow: initial one-time setup (orange), iterative annotation loop (green), and final review (gray).
Figure 2 presents the high-level workflow of BenchPress, which consists of a one-time setup phase followed by a repeated annotation loop. For non-nested SQL queries, the pipeline follows the standard sequence of steps, excluding the dashed steps. For nested queries, BenchPress automatically inserts two intermediate steps—decomposition and recomposition—to improve annotation accuracy and reduce complexity. These optional steps are depicted in the only dashed boxes in the diagram.
One-Time Setup:
  1. (1)
    Project Setup: The user selects or creates a new annotation project tied to a specific enterprise workload.
  2. (2)
    Data Ingestion: Users upload SQL logs and schema files or select from one of four supported public benchmarks: Bird, Fiben, Spider, or Beaver. The system parses and stores this data for further processing.
  3. (3)
    Task Configuration: Users choose the annotation direction (currently only SQL-to-NL), as well as the language model (e.g., GPT-4o, GPT-3.5 Turbo, or DeepSeek).
Annotation Loop:
  1. (3.5)
    (Optional) Decomposition: For nested SQL queries, the system rewrites the query into a series of Common Table Expressions (CTEs), breaking it down into logically independent subqueries.
  2. (4)
    Context Retrieval: For each SQL query or subquery, the system retrieves semantically similar examples using dense vector embeddings (e.g., Sentence-BERT Reimers and Gurevych (2019)). These examples consist of prior annotated queries (which naturally grow over time) and serve as guidance for generating relevant phrasing. In addition, BenchPress retrieves relevant tables with all their columns from the schema—either via SQL parsing (e.g., using sqlglot) or using the same embedding-based retrieval mechanism as for the examples. This combined context grounds the model’s output in both content and structure Lewis et al. (2020).
  3. (5)
    Candidate Generation: A large language model (LLM), selected in step 3, generates four candidate natural language descriptions for each SQL query or subquery. The prompt incorporates the retrieved examples and schema context from step 4 using a retrieval-augmented, few-shot prompting approach. Since examples can be long and dilute the prompt, the system always includes the relevant tables but only suggests the top-k retrieved examples to the user. This setup balances informativeness with prompt efficiency. We chose four candidates to balance linguistic diversity with annotation efficiency. Prior work in instruction tuning and human preference modeling often adopts this number as it provides sufficient variation while keeping the cognitive load manageable for human reviewers Ouyang et al. (2022); Wang et al. (2023). Generating multiple outputs also supports downstream use cases such as ranking, majority voting, or active learning.
  4. (5.5)
    (Optional) Recomposition: If decomposition was performed, the system automatically merges the subquery-level descriptions into a single coherent explanation of the original nested SQL query.
  5. (6)
    Feedback: Annotators can rank, refine, or discard priorities assigned to the LLM. This human-in-the-loop feedback improves prompt quality over time and supports future fine-tuning Christiano et al. (2017).
  6. (7)
    Review and Export: If ground truth annotations exist, outputs can be evaluated using automatic metrics (e.g., exact match, BLEU). Otherwise, users rely on qualitative assessment. Final annotations are exported in benchmark-ready format for training or evaluation.
The final exported annotations are then available in the typical benchmark format, i.e., JSON format, for downstream training and evaluation.
4.2. Key Design Features
In its design, BenchPress brings together several core techniques and strategies to support accurate and efficient annotation at scale:
  • Retrieval-Augmented Generation (RAG): To improve relevance, BenchPress retrieves semantically similar SQL queries and their annotations using dense vector search (e.g., via Sentence-BERT Reimers and Gurevych (2019)). These examples are embedded into prompts to ground the model in realistic phrasing patterns and schema usage Lewis et al. (2020). This approach aligns with best practices from the seminal Retrieval-Augmented Generation work Lewis et al. (2020), which shows significant accuracy gains for specialized NLP tasks through retrieval-enhanced prompting.
  • Prompt Engineering and Refinement: BenchPress utilizes structured prompt templates tailored explicitly to enterprise SQL logs. When initial suggestions fail to capture the intended meaning, annotators can iteratively refine prompts (e.g., emphasizing "filtering logic") to guide re-generation. This feedback-driven refinement loop has been shown to improve prompt quality and annotation accuracy, mirroring strategies from reinforcement learning from human feedback (RLHF) and human preference modeling Christiano et al. (2017).
  • Query Decomposition: For nested or structurally complex SQL queries, BenchPress decomposes them into simpler sub-queries. Natural language descriptions are then generated independently and later reassembled, reducing cognitive load and improving annotation precision.
  • Human-in-the-loop Feedback: At its core, BenchPress emphasizes continuous human oversight. Domain experts iteratively review, edit, rank, and flag unclear annotations produced by the model. This structured, iterative review phase ensures annotations meet enterprise-quality standards and reduces error propagation downstream. This reflects Google’s "PAIR" principles Dixon and Terry (2023) for responsible AI and supports findings that combining human judgment with AI outputs results in higher quality, robustness, and trustworthiness Amershi et al. (2019).
This modular architecture enables BenchPress to support diverse workflows and database schemas with minimal reconfiguration, while explicitly leveraging human expertise to maximize accuracy and adaptability for enterprise use cases.
5. User Study
5.1. Setup
To evaluate the impact of BenchPress on annotation efficiency and quality, we conducted a controlled user study using a between-subjects experimental design. In this design, each participant is randomly assigned to one condition only—ensuring that comparisons across conditions reflect differences in the interface or workflow, rather than learning effects or fatigue. This setup is widely used in HCI and behavioral research, including in experimental frameworks Lazar et al. (2017).
A total of 18 participants were recruited and first grouped into two strata—advanced and non-advanced SQL users—based on a pre-study questionnaire assessing their experience and familiarity with relational databases. Within each stratum, participants were randomly assigned to one of three experimental conditions using a balanced Latin square design to ensure counterbalancing:
  • Group A (BenchPress): Used the BenchPress interface, including schema information, example tables, logs, and four LLM-generated natural language suggestions per SQL query.
  • Group B (Manual): Provided only with schema files and logs, no LLM or suggestion support.
  • Group C (Vanilla LLM): Allowed to use a general-purpose LLM (e.g., ChatGPT) via its standard UI, but without RAG-based support or task-specific integration.
Each participant was assigned the same set of 30 SQL queries, sampled from the Beaver and Bird datasets (anonymized), and instructed to write a natural language description for each SQL query. The task was SQL-to-NL annotation only.
Independent variables in the study were the annotation condition (BenchPress, Manual, LLM) and user expertise (Advanced, Non-Advanced). The primary dependent variables included annotation time and annotation quality. Quality was assessed using both observational measures (e.g., back-translation match, ROUGE similarity) and participant self-reports on task confidence and perceived difficulty.
This setup enables a structured comparison of how different interfaces and user backgrounds affect performance in enterprise SQL annotation tasks.
5.2. Evaluation
We evaluated the performance of each annotation condition—BenchPress, Manual, and LLM—across three dimensions: annotation accuracy, annotation latency, and semantic fidelity using a backtranslation task.
Annotation Accuracy.
Annotation accuracy was measured by manually inspecting each NL description for fidelity to the corresponding SQL query. We checked whether key SQL components—such as column selections, calculations (e.g., aggregations), and grouping or ordering operations—were clearly and distinguishably described. As shown in Table 1, BenchPress consistently outperformed the Manual and LLM groups across both datasets. It produced more complete and structurally accurate descriptions, especially in complex enterprise queries from the Beaver dataset.
Avg Accuracy
BenchPress Vanilla LLM Manual Beaver 86.1% 66.2% 60.1% Bird 100.0% 100.0% 87.8% Overall 93.0% 83.1% 73.9%
Table 1. Annotation accuracy across BenchPress, Vanilla LLM, and Manual conditions on Beaver and Bird.
Annotation Latency.
Latency was measured as the total annotation time per participant, averaged across a given dataset. Table 2 shows that BenchPress led to the fastest annotation times, while the Manual group required by far the most time. This supports the hypothesis that context-aware LLM suggestions accelerate annotation without sacrificing quality.
Avg Latency
BenchPress Vanilla LLM Manual Beaver 16.1 min 16.2 min 102.1 min Bird 12.0 min 15.8 min 82.8 min Total 28.1 min 32.0 min 183.9min
Table 2. Average annotation latency (in minutes) per condition across all participants for each dataset.
Annotation Fidelity via Backtranslation.
To evaluate the semantic accuracy of the NL annotations produced in BenchPress, we performed a backtranslation task. In this process, an LLM was asked to regenerate SQL queries solely from the natural language descriptions created during annotation. The resulting SQL outputs were then compared to the original queries using a 5-level rubric.
We chose this rubric to capture both hard execution failures and finer semantic mismatches, providing a practical scale for measuring fidelity. Our rating system distinguishes between structural errors (e.g., incorrect tables or joins), content-level inaccuracies (e.g., wrong columns or filters), and minor deviations (e.g., ordering or phrasing issues). The five levels are defined as follows:
  • Level 1 – Invalid: The generated SQL fails to execute (e.g., due to syntax errors, undefined references, or broken nesting).
  • Level 2 – Executable but Structurally Incorrect: The SQL query runs but reflects major misunderstandings of the query’s structure. Examples include wrong tables, missing joins, or irrelevant subqueries.
  • Level 3 – Column-Level Errors: The SQL is structurally correct but uses incorrect columns, filters, functions, or groupings. The high-level intent is preserved, but the query’s logic is incorrect.
  • Level 4 – Minor Issues: The regenerated SQL is mostly faithful but contains small deviations such as incorrect sorting, missing nuance, or redundant clauses.
  • Level 5 – Fully Correct: The SQL matches the original in both structure and semantics, including all tables, conditions, filters, and ordering.
We used a vanilla LLM (without fine-tuning, chain-of-thought prompting, or in-context examples) to ensure that the results reflect the inherent information content of the NL descriptions, not artifacts of prompt engineering. This makes the evaluation a stricter test of how well the natural language alone communicates the SQL logic.
Backtranslation in this form provides a valuable lens on annotation quality: it captures not only whether a human would find a description understandable, but also whether it preserves enough detail for faithful round-tripping into executable SQL.
Figure 3. Backtranslation fidelity: proportion of SQL outputs at each clarity level across conditions.
Figure 3 shows that BenchPress yielded the highest proportion of 5 outputs, indicating superior semantic clarity. Manual and LLM groups more often fell into Level 4, typically due to row-order inconsistencies, omitted nuances, or subtle misinterpretations of intent.
5.3. Summary and Observations
Our user study yielded several notable insights:
Tool Effectiveness.
BenchPress consistently outperformed both the vanilla LLM and manual annotation approaches across all metrics—accuracy, efficiency, and semantic fidelity. These results demonstrate that integrating LLM-generated suggestions with lightweight user guidance significantly enhances the annotation process. The structured interface and contextual prompts provided by BenchPress not only improved output quality but also reduced cognitive load and task completion time.
Dataset Complexity.
We observed a clear divergence in tool performance between public and enterprise datasets. While all three tools performed reasonably well on the Bird dataset, the Beaver dataset—with its higher schema complexity, ambiguous column names, and enterprise-specific terminology—exposed substantial differences. BenchPress maintained high annotation quality in this more challenging setting, whereas the manual and vanilla LLM conditions struggled to preserve SQL semantics and coverage.
Implications for Benchmarking.
These findings highlight the limitations of existing tools and benchmarks when applied to real-world enterprise data. Public benchmarks do not adequately reflect the structural and linguistic complexity inherent in enterprise SQL logs. BenchPress addresses this gap by enabling scalable and accurate benchmark curation tailored to organizational data environments. By focusing on SQL-to-NL annotation with human-in-the-loop refinement, BenchPress provides a practical foundation for building robust, domain-specific text-to-SQL benchmarks and evaluating model performance in enterprise settings.
6. Conclusions and Future Work
BenchPress accelerates the creation of high-quality, domain-specific text-to-SQL benchmarks by combining LLM-based generation with human-in-the-loop validation. Our system uses retrieval-augmented prompting for intelligent SQL-to-text annotation and enables scalable evaluation of model performance on enterprise workloads. Through a controlled user study, we demonstrated that BenchPress improves annotation speed and fidelity, especially in typical enterprise settings with complex schemas and limited training data.
While the current system focuses on SQL-to-text annotation, a natural next step is to incorporate text-to-SQL generation for iterative validation. This would further increase the accuracy and speed of the benchmark curation process.
Another direction for future work is assessing the robustness of state-of-the-art models trained on public benchmarks such as Spider, Bird, and Fiben. Although many models achieve near-perfect performance on these datasets, it remains unclear whether they have overfit to canonical NL formulations. We plan to systematically rephrase the natural language queries in existing benchmarks—introducing more realistic, ambiguous, or underspecified variants—and re-evaluate model performance. This will reveal whether current benchmarks reflect genuine text-to-SQL generalization or merely reward surface-level pattern matching.
Creating custom benchmarks for domain-specific tasks with private data remains a significant bottleneck in the broader adoption of LLM applications. Data teams need better tools to rapidly create benchmarks that are representative of their unique data and use cases. BenchPress addresses this need by enabling adaptive, privacy-aware, and task-specific benchmarking of text-to-SQL systems within enterprise contexts.